A melting point test of n=10 samples of a binder used in manufacturing a rocket propellant resulted in x =154:2°F. Assume that the melting point is normally distributed with Q=1:5°F: Suppose you want to test the claim that the temperature is different from 155°F: Test the hypothesis at a=0:05: Write down the hypothesis to be tested and show all the steps and calculations.


1
Expert's answer
2021-07-07T17:32:35-0400

"H_0: \mu=155 \\

H_1: \mu ≠155 \\

α=0.05 \\

\bar{x}=154.2 \\

\sigma= 1.5 \\

n=10"

Test statistic:

"Z = \frac{\bar{x}- \mu}{\sigma / \sqrt{n}} \\

Z = \frac{154.2-155}{1.5/ \sqrt{10}}= -1.69 \\

Z_{α/2}=Z_{0.05/2} = 1.96"

Rejection criterion:

If "-Z_{\u03b1\/2}\u2264Z\u2264Z_{\u03b1\/2}" then we should fail to reject H0.

-1.69 > -1.96

"Z>Z_{\u03b1\/2}"

Accept H0.

The temperature is NOT different from 155°F